OpenStudy (anonymous):

5z^2=20z please help solve for z.

OpenStudy (saifoo.khan):

z = 4

OpenStudy (lalaly):

z^2-4z =0 z(z-4)=0 z=0 or z=4

OpenStudy (anonymous):

not only 4, but as lalaly wrote z=0 and z=4 . z^2-4z =0 z(z-4)=0 z=0 or z=4